# LiteTracker: Leveraging Temporal Causality for Accurate Low-latency Tissue Tracking

### v1

@@ Abstract
Tissue tracking plays a critical role in various surgical navigation and extended reality (XR) applications. While current methods trained on large synthetic datasets achieve high tracking accuracy and generalize well to endoscopic scenes, their runtime performances fail to meet the low-latency requirements necessary for real-time surgical applications. To address this limitation, we propose LiteTracker, a low-latency method for tissue tracking in endoscopic video streams. LiteTracker builds on a state-of-the-art long-term point tracking method, and introduces a set of training-free runtime optimizations. These optimizations enable online, frame-by-frame tracking by leveraging a temporal memory buffer for efficient feature re-use and utilizing prior motion for accurate track initialization. LiteTracker demonstrates significant runtime improvements being around 7×7$\times$7 × faster than its predecessor and 2×2$\times$2 × than the state-of-the-art. Beyond its primary focus on efficiency, LiteTracker delivers high-accuracy tracking and occlusion prediction, performing competitively on both the STIR and SuPer datasets. We believe LiteTracker is an important step toward low-latency tissue tracking for real-time surgical applications in the operating room.
Keywords:
Tissue Tracking Point Tracking Endoscopy.
@@ 1 Introduction
Figure 1: (Top) Demonstration of LiteTracker on a video from STIR Challenge 2024 dataset [16]. (Bottom) Latency and tracking accuracy comparison. Average Jaccard (AJ) metric is computed on SuPer dataset [11]. Latencies showcase the 95th percentile of inference step measurements for 1,024 points initalized at the first frame of a video with 615 frames. LiteTracker is approximately 2×2$\times$2 × faster than the fastest evaluated method, Track-On [1], and 7×7$\times$7 × than its predecessor, CoTracker3 [9], while exhibiting close to state-of-the-art tracking accuracy.
@@ 2 Methodology
We build LiteTracker setting CoTracker3 [9] as our foundation due to its high tracking accuracy and strong generalizibility to endoscopic scenes. In this section, we first give an overview of the CoTracker3 architecture and next, introduce our method, LiteTracker.
@@ 2.2.1 Frame-by-frame Processing with Temporal Memory
CoTracker3 [9] operates in sliding-window manner requiring accumulation of a stride ($S:=8$) of frames before performing the next prediction. A straight-forward way to perform frame-by-frame prediction is to set the stride $S$ to $1$ instead of $8$. This choice however, introduces a computational overhead since the processing of $T_{W}-S$ many frames need to be repeated multiple times. To tackle this challenge, we propose maintaining a temporal memory buffer to cache prior computations. Concretely, we cache the correlation features since their computation is relatively expensive, involving pair-wise similarity measurement of multi-scale patch features. To this end we maintain a ring buffer, with a capacity $T_{B}$ of 16 frames, adhering to the original window size of the CoTracker3. The ring buffer operates in a first-in, first-out (FIFO) manner where the newly processed frames overwrite the oldest ones. The temporal memory buffer is implemented in a way that it allows addition of new query points during the tracking process.
@@ 2.2.2 Track Initialization with Exponential Moving Average Flow
$\delta^{\text{avg}}$uparrow AJ $\uparrow$ OA $\uparrow$
CoTracker3 Online [9] Window $75.24$ $\mathbf{80.82}$ $\underline{96.96}$ $200.98$
Online BootsTAPIR [5] Frame $68.59$ $74.27$ $93.63$ $466.38$
MFT* [13] Frame $\mathbf{77.62}$ $67.02$ $83.73$ $317.05$
A-MFST* [2] Frame $58.59$
MFTIQ (SEA-RAFT)* [21] Frame $76.82$ $69.14$ $83.73$ $327.93$
MFTIQ (ROMA)* [21] Frame $\underline{77.22}$ $66.61$ $83.73$ $4355.11$
Track-On (Mem. size 12) [1] Frame $72.74$ $72.97$ $86.86$ $\underline{60.18}$
Track-On (Mem. size 48) [1] Frame $74.44$ $70.22$ $83.92$ $74.80$
LiteTracker (Ours) Frame $75.81$ $\underline{80.68}$ $\mathbf{97.45}$ $\mathbf{29.67}$
@@ 3.2.1 Datasets
We additionally evaluate LiteTracker on SuPer [11] dataset which consists of a single video of 522 frames, with manually annotated location and occlusion labels of 20 points for every $10^{\text{th}}$ frame. Leveraging the ground truth occlusion labels, we employ the average Jaccard (AJ) and occlusion accuracy (OA) metrics to measure the tracking and visbility prediction performance. Following their original implementation [4], we compute AJ on the image size of $256\times 256$ with accuracy tresholds set to $\{1,2,4,8,16\}$.
@@ 3.2.2 Results
Figure 3: Qualitative results on video samples from STIR Challenge 2024 [16] (top) and StereoMIS [7] (bottom) datasets. LiteTracker shows high tissue-tracking accuracy and occlusion handling under challenging deformations, tool interactions and perspective changes.
@@ 3.2.3 Ablation Studies
In our first study, we evaluate EMA flow intialization’s contribution to the overall tracking accuracy with respect to the varying number of refinement iterations. Our results, collected on the STIR Challenge 2024 dataset [16], showcase that our EMA flow initialization already yields the model’s highest tracking accuracy with a single pass through the iterative refinement module. We believe that our initialization scheme mimics temporal smoothness common to deformations in surgical scenes, there by provides a robust first estimate. Unlike CoTracker3, see Sec.3.1, further refinement iterations degrade the tracking accuracy of LiteTracker. We believe this can be attributed to the use of pre-trained weights.
@@ 4 Conclusion
In this paper, we propose LiteTracker, a low-latency tissue tracking method designed for real-time surgical applications. Building on CoTracker3, a state-of-the-art long-term point tracking method, LiteTracker incorporates a set of training-free runtime optimizations enabling efficient and accurate frame-by-frame tissue tracking. Our experimental results demonstrate that LiteTracker achieves substantial runtime improvements, running approximately 7×7$\times$7 × faster than its predecessor and 2×2$\times$2 × faster than the state-of-the-art, making it well-suited for real-time deployment. Furthermore, despite being designed for efficiency, LiteTracker maintains competitive tracking accuracy, coming close to the state-of-the-art on the SuPer and STIR datasets. We believe LiteTracker presents a significant step toward practical usage of low-latency, accurate tissue tracking and opens new avenues for future research utilizing it as a building block for real-time surgical applications.

### v2

@@ Abstract
Tissue tracking plays a critical role in various surgical navigation and extended reality (XR) applications. While current methods trained on large synthetic datasets achieve high tracking accuracy and generalize well to endoscopic scenes, their runtime performances fail to meet the low-latency requirements necessary for real-time surgical applications. To address this limitation, we propose LiteTracker, a low-latency method for tissue tracking in endoscopic video streams. LiteTracker builds on a state-of-the-art long-term point tracking method, and introduces a set of training-free runtime optimizations. These optimizations enable online, frame-by-frame tracking by leveraging a temporal memory buffer for efficient feature re-use and utilizing prior motion for accurate track initialization. LiteTracker demonstrates significant runtime improvements being around $7\times$ faster than its predecessor and $2\times$ than the state-of-the-art. Beyond its primary focus on efficiency, LiteTracker delivers high-accuracy tracking and occlusion prediction, performing competitively on both the STIR and SuPer datasets. We believe LiteTracker is an important step toward low-latency tissue tracking for real-time surgical applications in the operating room. Our code is publicly available at https://github.com/ImFusionGmbH/lite-tracker.
@@ 1 Introduction
Figure 1: (Top) Demonstration of LiteTracker on a video from STIR Challenge 2024 dataset [16]. (Bottom) Latency and tracking accuracy comparison. Average Jaccard (AJ) metric is computed on SuPer dataset [11]. Latencies showcase the 95th percentile of inference step measurements for 1,024 points initalized at the first frame of a video with 615 frames. LiteTracker is approximately $2\times$ faster than the fastest evaluated method, Track-On [1], and $7\times$ than its predecessor, CoTracker3 [9], while exhibiting close to state-of-the-art tracking accuracy.
@@ 2 Methodology
We build LiteTracker using CoTracker3 [9] as our foundation due to its high tracking accuracy and strong generalizibility to endoscopic scenes. In this section, we first give an overview of the CoTracker3 architecture and next, introduce our method, LiteTracker.
@@ 2.2.1 Frame-by-frame Processing with Temporal Memory
CoTracker3 [9] operates in a sliding-window manner requiring accumulation of a stride ($S:=8$) of frames before performing the next prediction. A straight-forward way to perform frame-by-frame prediction is to set the stride $S$ to $1$ instead of $8$. This choice however, introduces a computational overhead since the processing of $T_{W}-S$ many frames need to be repeated multiple times. To tackle this challenge, we propose maintaining a temporal memory buffer to cache prior computations. Concretely, we cache the correlation features since their computation is relatively expensive, involving pair-wise similarity measurement of multi-scale patch features. To this end we maintain a ring buffer, with a capacity $T_{B}$ of 16 frames, adhering to the original window size of CoTracker3. The ring buffer operates in a first-in, first-out (FIFO) manner where the newly processed frames overwrite the oldest ones. Our temporal memory buffer is expressly designed to allow the introduction of new query points at any time during tracking. When additional points are queried in later frames, a set of proxy feature vectors are appended to the buffer. The transformer architecture dynamically masks these proxies during temporal and spatial attention, ensuring that they do not influence existing tokens while preserving the causality of the pipeline.
@@ 2.2.2 Track Initialization with Exponential Moving Average Flow
$\delta^{\text{avg}}$ $\uparrow$ AJ $\uparrow$ OA $\uparrow$
A-MFST* Frame 58.59
MFTIQ (ROMA)* Frame 77.22 66.61 83.73 4355.11
Online BootsTAPIR Frame 68.59 74.27 93.63 466.38
MFTIQ (SEA-RAFT)* Frame 76.82 69.14 83.73 327.93
MFT* Frame 77.62 67.02 83.73 317.05
CoTracker3 Online Win. 75.24 80.82 96.96 200.98
Track-On (48) Frame 74.44 70.22 83.92 74.80
Track-On (12) Frame 72.74 72.97 86.86 60.18
LiteTracker (Ours) Frame 75.81 80.68 97.45 29.67
@@ 3.2.1 Datasets
We additionally evaluate LiteTracker on the SuPer [11] dataset which consists of a single video of 522 frames, with manually annotated location and occlusion labels of 20 points for every $10^{\text{th}}$ frame. Leveraging the ground truth occlusion labels, we employ the average Jaccard (AJ) and occlusion accuracy (OA) metrics to measure the tracking and visibility prediction performance. Following their original implementation [4], we compute AJ on the image size of $256\times 256$ with accuracy thresholds set to $\{1,2,4,8,16\}$.
@@ 3.2.2 Results
Figure 3: Qualitative results on video samples from the STIR Challenge 2024 [16] (top) and StereoMIS [7] (bottom) datasets. LiteTracker shows high tissue-tracking accuracy and occlusion handling under challenging deformations, tool interactions and perspective changes.
@@ 3.2.3 Ablation Studies
In our first study, we evaluate the EMA flow initialization’s contribution to the overall tracking accuracy with respect to the varying number of refinement iterations. Our results, collected on the STIR Challenge 2024 dataset [16], showcase that our EMA flow initialization already yields the model’s highest tracking accuracy with a single pass through the iterative refinement module. We believe that our initialization scheme mimics temporal smoothness common to deformations in surgical scenes, there by providing a robust first estimate. Unlike CoTracker3, see Sec.3.1, further refinement iterations degrade the tracking accuracy of LiteTracker. We believe this can be attributed to the use of pre-trained weights.
@@ 4 Conclusion
In this paper, we propose LiteTracker, a low-latency tissue tracking method designed for real-time surgical applications. Building on CoTracker3, a state-of-the-art long-term point tracking method, LiteTracker incorporates a set of training-free runtime optimizations enabling efficient and accurate frame-by-frame tissue tracking. Our experimental results demonstrate that LiteTracker achieves substantial runtime improvements, running approximately $7\times$ faster than its predecessor and $2\times$ faster than the state-of-the-art, making it well-suited for real-time deployment. Furthermore, despite being designed for efficiency, LiteTracker maintains competitive tracking accuracy, coming close to the state-of-the-art on the SuPer and STIR datasets. We believe LiteTracker presents a significant step toward practical usage of low-latency, accurate tissue tracking and opens new avenues for future research utilizing it as a building block for real-time surgical applications.